# Are LLMs Smarter Than Chimpanzees? An Evaluation on Perspective Taking and Knowledge State Estimation

### v1

@@ Abstract
Cognitive anthropology suggests that the distinction of human intelligence lies in the ability to infer other individuals’ knowledge states and understand their intentions Tomasello et al. (2005). In comparison, our closest animal relative, chimpanzees, lack the capacity to do so Call and Tomasello (2008); Hare (2011). With this paper, we aim to evaluate LLM performance in the area of knowledge state tracking and estimation. We design two tasks to test (1) if LLMs can detect when story characters, through their actions, demonstrate knowledge they should not possess, and (2) if LLMs can predict story characters’ next actions based on their own knowledge vs. objective truths they do not know. Results reveal that most current state-of-the-art LLMs achieve near-random performance on both tasks, and are substantially inferior to humans. We argue future LLM research should place more weight on the abilities of knowledge estimation and intention understanding.
@@ 1 Introduction
Figure 1: The Knowledge-sensitive Next-action Prediction (KNP) task, where the implausible knowledge is highlighted for easy reading (but not highlighted for the LLMs and the human test-takers). This task tests whether LLMs can take the perspective of the story characters and choose actions based on knowledge they can access.
@@
Dataset # Samples Context Length Random Chance GPT-4o Accuracy
Long-context Benchmarks
Needle Threading Roberts et al. (2024) 660 1.2K tokens $\approx 0$ 0.93
BABILong (QA1) Kuratov et al. (2024) 100 2.0K tokens $\approx 0$ 0.94
Theory-of-mind and Narrative Benchmarks
FANTOM Kim et al. (2023) 10K 537 words $\approx 0$ 0.80
SIMPLETOM (BehaviorQA) Gu et al. (2024) 1.1K 38 words 0.50 0.94
CharTOM-QA (BeliefQA) Zhou et al. (2025) 201 2K words 0.25 0.58
FlawedFictions ClssificationAhuja et al. (2025) 414 731 words 0.50 0.64
@@
IKD Classification (Ours) 1K 588 words 0.50 0.50
KNP (Ours) 500 98 words 0.50 0.51
Table 1: Comparison between several datasets relying on long-context understanding, theory-of-mind, and narrative understanding, and the performance of GPT-4o. Even though each story of the proposed benchmarks is short, they create harder challenges for GPT-4o than existing benchmarks.
@@
Therefore, a natural question to ask is: can LLMs infer the knowledge states of human actors? Using textual stories, we design two probing tasks: (1) Implausible Knowledge Detection (IKD) (Fig. 2), which tests whether LLMs can detect when a story character demonstrates knowledge they should not possess, and (2) Knowledge-sensitive Next-action Prediction (KNP) (Fig. 1), which tests whether LLMs can predict actions that story characters will likely take, depending on their knowledge.
As shown in Figure 2, we construct two benchmark datasets, one for each task. For IKD, we collect stories from the internet111www.reedsy.com and www.short-story.me, summarize them with an LLM, and rewrite the story to introduce a plot hole Ahuja et al. (2025). In the plot hole, a story character uses knowledge they cannot possibly possess in their action. The LLM is required to detect if a plot hole exists in a given story. The benchmark consists of 500 correct stories and 500 erroneous stories in 5 genres.
For the KNP task, we rewrite the stories with an LLM such that each story contains a knowledge statement that cannot be plausibly known by a story character, but has important implications on their next action. The LLM being tested must predict the next action of the story character, and the correct choice is that the character will act without the knowledge. The dataset consists of 500 QA pairs. Both IKD and KNP stories were validated by human annotators. Section 4 contains more details.
Despite their apparent simplicity (see comparisons with other datasets in Table 1), the proposed benchmarks turn out to be highly adversarial to state-of-the-art LLMs. On the KNP task, humans with zero training and zero demonstrations achieve an accuracy of 92%, 21 percentage points (pp) above the best LLM, Claude-4.5-Opus. The best GPT-5 model achieves only 63%, whereas GPT-4o performs at random chance. On the task of IKD, Claude-4.5-Opus achieves 66%. All other models perform below 60%, while random chance is 50%. The zero-shot human baseline is at 76%. The surprising failures of LLMs in understanding seemingly simple stories highlight their deficiency in the critical ability of estimating knowledge states.
@@
  • Inspired by cognitive research, we propose an important dimension of LLM evaluation: the ability to understand and share intentions (USI) of humans. As USI can be difficult to measure directly, we propose to evaluate a prerequisite skill, knowledge state tracking and estimation (KSTE).
  • We build simple and short KSTE tests that push most state-of-the-art LLMs (except the latest Claude and Gemini models) to marginally above random chance. The finding is interesting by itself and highlights the need for further research on USI.
@@ 2 Related Works
Fundamental to social intelligence, Theory of Mind refers to the ability to recognize the fact that other people possess distinct mental states and to estimate such states Nguyen et al. (2025). Most studies that test ToM in LLMs focus on the dimension of beliefs and knowledge Kosinski (2024); Kim et al. (2023); Wu et al. (2023); Nematzadeh et al. (2018); Xu et al. (2024), others cover intentions Staab et al. (2023); Zhou et al. (2023) and emotions Wu et al. (2024).
In this work, we focus on the dimension of knowledge. Existing works face two limitations: 1) Most follow the false-belief test Baron-Cohen et al. (1985), which tests if the LLM can infer whether a person knows about a change in the world (e.g., a milk bottle having been moved from the fridge to the table). The common problem setting provides two alternative world states for LLMs to choose from, making the problem easier. Our approach, in contrast, presents one fact that the character should not know about. Without the alternative choice, the LLM may inadvertently consider the fact as the only possible knowledge state. 2) The questions explicitly ask about story character’s knowledge state (e.g., “Where does Alice believe the milk is?”). This gives LLMs a hint to perform knowledge tracking. In contrast, testing the LLMs on the action implications of knowledge (our setting) is harder, using questions like “Alice wants to drink milk. What would she do? A) Open the fridge. B) Go to the table.” The inference that the knowledge state is critical in this decision appears to be rather difficult for LLMs (see the first ablation in §5.2) but is vital for USI.
Additionally, compared to synthetic datasets that rely on standardized templates, such as those based on object locations Wu et al. (2023); Xu et al. (2024), free-form narratives offer more diverse and potentially out-of-distribution assessments for ToM Van Duijn et al. (2023). CharToM-QA Zhou et al. (2025) uses classic novels to explore ToM understanding, covering different characters’ perspectives. However, their questions are based on lengthy novels (over 2K words), making it unclear whether the challenge stems from context length or ToM understanding. Our dataset, by contrast, features much shorter natural narratives (Figure 1), yet LLMs still show near-random performance.
@@ 3 Tasks of Implausible Knowledge
We design two tasks to evaluate LLMs’ understanding of knowledge. The IKD task (Sec 3.1) focuses on the detection of actions that depend on implausible knowledge ex post, whereas the KNP task (Sec 3.2) focuses on the prediction of next actions ex ante. As such, these two tasks test two complementary aspects of the understanding of implausible knowledge.
3.1 Task 1: Implausible Knowledge Detection
@@
The task further contains two subtasks. In the binary classification subtask, the LLM classifies each story into logically consistent (1) or inconsistent (0). We report the classification accuracy. Given a story $x_{i}$, a predicted class $\hat{y}_{i}$, and the ground-truth label $y_{i}$, the accuracy is simply $\sum_{i}^{N}\mathbbm{1}(\hat{y}_{i}=y_{i})/N$, where $N$ is the number of stories and $\mathbbm{1}(\cdot)$ is the indicator function.
@@
Figure 2: Our dataset construction process (Section 4). Using the original story, we first identify implausible knowledge that a character cannot possess at a given event, and that would affect the character’s action. If no such knowledge exists, we prompt the model to introduce one. Then: 1) For the IKD task, we rewrite the original event to generate an erroneous story; 2) For the KNP task, we construct a question about a knowledge-sensitive action, creating two answer options—one without the implausible knowledge and one with the knowledge.
3.2 Task 2: Knowledge-sensitive Next-action Prediction
@@ 4.2 Data Construction
Combined with the 500 erroneous and 500 original correct stories [^3] Original stories are also verified and refined by human annotators to remove any possible logic errors., this yields the dataset for the IKD task, with an average length of 588 words.
Next-action Question Construction.
Human annotators check for the correctness of the answers. For example, suppose $\mathcal{K}$ is “the girl whom $C$ loves thinks he is weird,” and $C$ is characterized as very shy. If the correct option is “he finds the courage to confess” and the distractor is “he does not have the nerve to confess”, the distractor may remain plausible solely due to shyness rather than being caused by $\mathcal{K}$. In this case, human annotators may choose to rewrite the options themselves or regenerate using the LLM. The final refined dataset contains 500 QA pairs for the KNP task, with stories averaging 98 words in length.
@@
Implausible Knowledge Detection KNP
Binary Classification Localization Action Choice
Pos Cls Neg Cls Full Cls Neg Loc Full Loc Accuracy
@@
Qwen3-8B 0.59 0.31 0.45 0.04 0.31 0.57
Qwen3-8B (think) 0.64 0.40 0.52 0.08 0.36 0.56
Qwen3-32B 0.67 0.36 0.51 0.07 0.37 0.52
Qwen3-32B (think) 0.60 0.49 0.55 0.07 0.33 0.55
Qwen3-Max 0.75 0.36 0.56 0.13 0.44 0.60
LLaMA3.1-8B 0.19 0.68 0.44 0.11 0.15 0.51
LLaMA3.1-70B 0.47 0.56 0.52 0.08 0.27 0.54
LLaMA3.3-70B 0.71 0.30 0.51 0.05 0.37 0.53
Kimi-K2-Instruct 0.29 0.73 0.51 0.19 0.24 0.55
Kimi-K2-Thinking 0.19 0.87 0.53 0.37 0.28 0.56
DeepSeek-V3.1 0.41 0.70 0.56 0.22 0.32 0.53
DeepSeek-R1-V3.1 0.70 0.47 0.58 0.23 0.46 0.58
DeepSeek-V3.2 0.62 0.56 0.58 0.22 0.42 0.55
DeepSeek-R1-V3.2 0.89 0.29 0.59 0.19 0.54 0.57
@@
GPT-4o 0.91 0.09 0.50 0.03 0.47 0.51
GPT-5-mini 0.74 0.22 0.48 0.06 0.40 0.54
GPT-5 (low) 0.73 0.41 0.57 0.17 0.45 0.62
GPT-5 (medium) 0.68 0.41 0.55 0.16 0.42 0.63
GPT-5 (high) 0.75 0.38 0.57 0.14 0.45 0.63
Gemini-3-Pro 0.23 0.92 0.56 0.48 0.36 0.70
Claude-4.5-Opus 0.80 0.51 0.66 0.31 0.56 0.71
Claude-4.5-Sonnet 0.29 0.86 0.58 0.34 0.32 0.67
Zero-shot Human 0.82 0.70 0.76 0.61 0.72 0.92
Table 2: Main results on IKD and KNP tasks. For IKD, we show the classification and localization performance on the positive set (correct stories), negative set (erroneous stories), and the full set.
@@ 5 Experiments
We test several models across two categories: closed-source models (Gemini-3-Pro, GPT-5, GPT-5-mini, GPT-4o, Claude-4.5-Sonnet, and Claude-4.5-Opus) and open-source models (LLaMA3.3-70B Dubey et al. (2024a), LLaMA3.1-70B Dubey et al. (2024b), LLaMA3.1-8B, Qwen3-MAX, Qwen3-32B Yang et al. (2025), Qwen3-8B, DeepSeek-Chat, DeepSeek-R1 (both v3.1 and v3.2) Liu et al. (2025), Kimi-K2-Instruct Team et al. (2025), and Kimi-K2-Thinking.
@@
We test the LLMs using the zero-shot setting. Since the task requires reasoning about knowledge states, we always ask the models to generate their thinking process and output the answer. For temperature and top-p settings, we apply the default settings for all open-source models and closed-source APIs. For the KNP task, we randomly shuffle the position of correct and distractor answers to avoid positional bias in LLMs.
@@ 5.1 Main Results
As shown in Table 2, most models achieve near-random performance. For logic error detection (IKD), many leading models perform only slightly better than random (0.5) in the binary classification task. For instance, GPT-5 (high) achieves 0.57, DeepSeek-R1 reaches 0.59, and Claude-4.5-Opus shows the highest performance at 0.66. In the localization task, performance is even lower—DeepSeek-R1 achieves 0.54, and Gemini-3-Pro reaches 0.36. An interesting finding is that different models also exhibit distinct judgment tendencies: GPT-4o shows a strong bias toward determining stories are logically consistent, while Gemini-3-Pro shows the opposite tendency. As a result, although some models display high performance on either the positive set (stories with no errors) or the negative set, the overall performance is poor. Human performance is also not particularly high, largely due to the lengthy context. If we highlight the events containing the error and the relevant knowledge, their localization performance on negative set increases from 0.61 to 0.86.
For knowledge-sensitive action prediction (KNP), most models also perform near-randomly — especially open-source models. Among leading models, Gemini-3-Pro and Claude-4.5-Opus achieve the highest performance at around 0.7. Analysis of their reasoning processes reveals that they consider knowledge states when predicting character actions (Case shown in Fig. 10). However, their understanding of knowledge states still needs improvement, revealing a gap with human performance of 0.92.
@@
Takeaway Most state-of-the-art LLMs show near-random performance in knowledge estimation and tracking. Chain-of-thought reasoning does not show clear advantage.

### v2

@@ Abstract
Cognitive anthropology suggests that the distinction of human intelligence lies in the ability to infer other individuals’ knowledge states and understand their intentions Tomasello et al. (2005). In comparison, our closest animal relative, chimpanzees, lack the capacity to do so Call and Tomasello (2008); Hare (2011). With this paper, we aim to evaluate LLM performance in estimating other individuals’ knowledge states and their potential actions. We design two tasks to test (1) if LLMs can predict story characters’ next actions based on their own knowledge vs. improperly using information unavailable from their perspective, and (2) if LLMs can detect when story characters, through their actions, demonstrate knowledge they should not possess. Results reveal that most current state-of-the-art LLMs achieve near-random performance on both tasks, and are substantially inferior to humans. We argue future LLM research should place more weight on the abilities of knowledge estimation and intention understanding. The project is available at: https://github.com/DingyiYang/Act-on-Known/.
@@ 1 Introduction
Figure 1: The Knowledge-sensitive Next-action Prediction (KNP) task, where the implausible knowledge is highlighted for easy reading (but not highlighted for the LLMs and the human test-takers). This tests whether LLMs can take the perspective of the story characters and choose actions based on knowledge they can access.
@@
Ability Tested Knowledge Domain #Samples Story Length Performance
Knowledge State Tracking Potential Action Estimation In-Context Localization Knowledge Domain Knowledge Creation Random GPT-5.4 (high)
Theory-of-Mind Benchmarks
ToMi Le et al. (2019) - - Location Rule-Based 6K 40 words 0.50 0.90
FANTOM Kim et al. (2023) - - 3 Types LLM 10K 537 words 0.50 0.76
T4D Zhou et al. (2023a) - Location LLM - - - -
SIMPLETOM (BehaviorQA) Gu et al. (2026) - 10 Types LLM 1.1K 38 words 0.50 0.77
CharTOM-QA (BeliefQA) Zhou et al. (2025) - General Classic Novel 201 2K words 0.25 0.65
@@
KNP (Ours) - General Human & LLM 500 93 words 0.50 0.61
IKD (Ours) General Human & LLM 1K 601 words 0.25 0.53
Table 1: Comparison between existing datasets. KNP requires models to track characters’ knowledge states and estimate their potential next actions. IKD requires models to do so for each event in context, and to localize any implausible action. Compared to datasets that rely on limited LLM-generated scenarios or classic novels prone to data leakage, our datasets cover broader situations and require richer commonsense understanding of the real world.
@@
Therefore, a natural question to ask is: can LLMs infer the knowledge states of human actors? Using textual stories, we propose two probing tasks: (1) Knowledge-sensitive Next-action Prediction (KNP) (Fig. 1), which tests whether LLMs can predict actions that story characters will likely take, depending on their own knowledge. (2) Implausible Knowledge Detection (IKD) (Fig. 2), which tests whether LLMs can detect when a story character acts on knowledge they should not possess.
As shown in Figure 2, we construct two benchmark datasets, one for each task. We collect and summarize human-written stories from the internet111www.reedsy.com and www.short-story.me. For KNP, we identify or create  [^2] If no such knowledge exists in the original story, we use LLM to introduce one. a knowledge statement that cannot be plausibly known by a story character, but has important implications for their next action. The LLM being tested must predict the next action of the character, and the correct choice is that the character will act without the knowledge. The dataset consists of 500 QA pairs in 5 genres.
For the IKD task, we rewrite the story to introduce a plot hole Ahuja et al. (2025). In the plot hole, a story character uses knowledge they cannot possibly possess in their action. The LLM is required to detect if a plot hole exists in a given story. The benchmark consists of 500 correct stories and 500 erroneous stories. Both KNP and IKD stories were validated by human annotators. Section 4 contains more details.
Our proposed benchmarks demand a more advanced capacity for tracking knowledge states and understanding how they affect actions in the real world (Table 1 shows comparisons with existing datasets). They are highly adversarial to state-of-the-art LLMs. On the KNP task, humans with zero training and zero demonstrations achieve an accuracy of 92%, 23 percentage points (pp) above the best LLM, Claude-4.5-Opus. GPT-5.4 achieves only 61%, whereas GPT-4o performs at random chance. On the IKD task, Claude-4.5-Opus achieves 68% classification accuracy and 0.59 IoU for localization, whereas other models remain at or below roughly 60% accuracy. The zero-shot human baseline reaches 76% accuracy. The surprising failures of LLMs in understanding seemingly simple stories highlight their deficiency in the critical ability of estimating knowledge states and potential actions.
@@
  • Inspired by cognitive research, we propose an important dimension of LLM evaluation: the ability to understand and share intentions (USI) of humans. As USI can be difficult to measure directly, we propose to evaluate a prerequisite skill: knowledge state tracking and potential action estimation.
  • We build simple and short tests that push most state-of-the-art LLMs (except the latest Claude and Gemini models) to marginally above random chance. The finding is interesting by itself and highlights the need for further research on USI.
@@ 2 Related Works
Fundamental to social intelligence, Theory of Mind refers to the ability to recognize the fact that other people possess distinct mental states and to estimate such states Nguyen and others (2025). Most studies that test ToM in LLMs focus on the dimension of beliefs and knowledge Kosinski (2024); Kim et al. (2023); Wu et al. (2023); Nematzadeh et al. (2018); Xu et al. (2024), others cover intentions Staab et al. (2023); Zhou et al. (2023b) and emotions Wu et al. (2024).
In this work, we focus on the dimension of knowledge. Most existing works face two limitations: 1) they explicitly ask about knowledge state understanding, like in ToMi Le et al. (2019), Hi-ToM Wu et al. (2023), and FANTOM Kim et al. (2023). For example, given a story “Sally is in one room; in another room, Anne puts milk in a box,” they ask “What does Sally think is in the box?”, which directly cues LLMs to perform knowledge tracking. In contrast, testing the LLMs on the action implications of knowledge (our setting) is harder, using questions like “Sally wants to drink milk. What would she do? A: Go to the box in another room. B: Try to find milk.” The inference that the knowledge state is critical in this decision appears to be rather difficult for LLMs (see the first ablation in §5.2) but is vital for USI. 2) Many works employ false-belief contexts, where the story presents a world state change (e.g., “Sally puts milk in a basket and leaves; Anne moves it to a box”), like in T4D Zhou et al. (2023a) and Open-TOM Xu et al. (2024). This setup provides two alternative world states, implicitly letting LLMs choose between them. In contrast, our stories present only an unchanging fact that the character cannot know. If LLMs fail to understand that one’s action is based on their own knowledge, they will directly use this inaccessible fact. This design more rigorously tests whether LLMs understand how perspective-specific knowledge affects action.
Additionally, compared to datasets built from standardized templates Le et al. (2019) or limited scenarios Gandhi et al. (2023); Chen et al. (2024); Gu et al. (2026), our dataset is drawn from diverse human-written stories, which cover a broader range of realistic scenarios and offer a more out-of-distribution assessment for ToM. Although CharToM-QA Zhou et al. (2025) also targets general scenarios, it relies on classic novels and is subject to serious data leakage concerns. Moreover, the lengthy context (over 2K words) makes it unclear whether the challenge stems from context length or ToM understanding. Our KNP task, by contrast, features much shorter natural narratives (93 words on average), yet LLMs still show near-random performance.
@@ 3 Tasks of Implausible Knowledge
We design two tasks to evaluate LLMs’ understanding of knowledge. The KNP task (Sec 3.1) focuses on the prediction of next actions ex ante, whereas the IKD task (Sec 3.2) focuses on the detection of actions that depend on implausible knowledge ex post. As such, these two tasks test two complementary aspects of the understanding of implausible knowledge.
3.1 Task 1: Knowledge-sensitive Next-action Prediction
3.2 Task 2: Implausible Knowledge Detection
@@
The task contains two subtasks. In the binary classification subtask, the LLM classifies each story into logically consistent (1) or inconsistent (0). We report the classification accuracy. Given a story $x_{i}$, a predicted class $\hat{y}_{i}$, and the ground-truth label $y_{i}$, the accuracy is simply $\sum_{i}^{N}\mathbbm{1}(\hat{y}_{i}=y_{i})/N$, where $N$ is the number of stories and $\mathbbm{1}(\cdot)$ is the indicator function.
@@
Figure 2: Our dataset construction process (Section 4). Using the original story, we first identify implausible knowledge that a character cannot possess at a given event, and that would affect the character’s action. If no such knowledge exists, we prompt the model to introduce one. Then: 1) For the KNP task, we construct a question about a knowledge-sensitive action, creating two answer options—one without the implausible knowledge and one with the knowledge; 2) For the IKD task, we rewrite the original event to generate an erroneous story.
@@ 4.2 Data Construction
Next-action Question Construction.
Human annotators check for the correctness of the answers. For example, suppose $\mathcal{K}$ is “the girl whom $C$ loves thinks he is weird,” and $C$ is characterized as very shy. If the correct option is “he finds the courage to confess” and the distractor is “he does not have the nerve to confess”, the distractor may remain plausible solely due to shyness rather than being caused by $\mathcal{K}$. In this case, human annotators may choose to rewrite the options themselves or regenerate using the LLM. The final refined dataset contains 500 QA pairs for the KNP task, with stories averaging 93 words in length.
@@
Combining the 500 erroneous and 500 original correct stories [^4] Original stories are also verified and refined by human annotators to remove any possible logic errors., this yields the dataset for the IKD task, with an average length of 601 words.
@@
KNP Implausible Knowledge Detection (IKD)
Action Choice Binary Classification Localization
Accuracy Pos Cls Neg Cls Full Cls Neg Loc Full Loc
@@
Qwen3-8B 0.50 0.58 0.31 0.45 0.04 0.31
Qwen3-8B (think) 0.49 0.69 0.32 0.51 0.05 0.37
Qwen3-32B 0.49 0.78 0.23 0.51 0.04 0.41
Qwen3-32B (think) 0.50 0.65 0.45 0.55 0.10 0.38
Qwen3-Max (think) 0.53 0.86 0.24 0.55 0.07 0.47
LLaMA3.1-8B 0.51 0.14 0.64 0.39 0.11 0.13
LLaMA3.1-70B 0.55 0.41 0.60 0.51 0.08 0.25
LLaMA3.3-70B 0.54 0.69 0.36 0.53 0.05 0.37
Kimi-K2.5-Instruct 0.56 0.71 0.51 0.61 0.27 0.49
Kimi-K2.5-Thinking 0.57 0.74 0.49 0.62 0.26 0.50
DeepSeek-V3.1 0.47 0.40 0.65 0.53 0.16 0.28
DeepSeek-R1-V3.1 0.55 0.76 0.46 0.61 0.22 0.49
DeepSeek-V3.2 0.50 0.68 0.50 0.59 0.20 0.44
DeepSeek-R1-V3.2 0.57 0.94 0.27 0.61 0.16 0.55
@@
GPT-4o 0.49 0.92 0.10 0.51 0.03 0.48
GPT-5-mini 0.52 0.83 0.18 0.51 0.06 0.45
GPT-5 (low) 0.57 0.81 0.37 0.59 0.18 0.50
GPT-5 (medium) 0.58 0.80 0.33 0.57 0.15 0.48
GPT-5 (high) 0.58 0.85 0.30 0.58 0.13 0.49
GPT-5.4 (high) 0.61 0.93 0.27 0.60 0.12 0.53
Gemini-3-Pro 0.67 0.28 0.86 0.57 0.47 0.38
Claude-4.5-Sonnet 0.66 0.35 0.82 0.59 0.29 0.32
Claude-4.5-Opus 0.69 0.88 0.48 0.68 0.30 0.59
Zero-shot Human 0.92 0.82 0.70 0.76 0.61 0.72
Table 2: Main results on KNP and IKD tasks. For IKD, we show the classification and localization performance on the positive set (correct stories), negative set (erroneous stories), and the full set.
@@ 5 Experiments
We test several models across two categories: closed-source models (Gemini-3-Pro, GPT-5.4, GPT-5, GPT-5-mini, GPT-4o, Claude-4.5-Sonnet, and Claude-4.5-Opus) and open-source models (LLaMA3.3-70B Dubey et al. (2024b), LLaMA3.1-70B Dubey et al. (2024a), LLaMA3.1-8B, Qwen3-MAX, Qwen3-32B Yang et al. (2025), Qwen3-8B, DeepSeek-Chat, DeepSeek-R1 (v3.1 and v3.2) Liu et al. (2025), Kimi-K2.5-Instruct Team et al. (2026), and Kimi-K2.5-Thinking.
@@
We test the LLMs using the zero-shot setting. Since the task requires reasoning about knowledge states, we always ask the models to generate their thinking process and output the answer. For temperature and top-p settings, we apply the default settings for all open-source models and closed-source APIs, then average scores across 3 runs. For the KNP task, we randomly shuffle the position of correct and distractor answers to avoid positional bias in LLMs.
@@ 5.1 Main Results
The results are shown in Table 2. For knowledge-sensitive action prediction (KNP), most models perform near-randomly — especially open-source models. Among leading models, Claude-4.5-Opus and Gemini-3-Pro achieve the highest performance at around 0.7. Analysis of their reasoning processes reveals that they consider knowledge states when predicting character actions (Case shown in Fig. 10). However, their understanding of knowledge states still needs improvement, revealing a gap with human performance of 0.92.
For logic error detection (IKD), many leading models perform only slightly better than random in the binary classification task. For instance, GPT-5 (high) achieves 0.58, DeepSeek-R1 reaches 0.61, and Claude-4.5-Opus shows the highest performance at 0.68. In the localization task, performance is even lower—DeepSeek-R1 achieves 0.55, and Gemini-3-Pro reaches 0.38. An interesting finding is that different models also exhibit distinct judgment tendencies: GPT-4o shows a strong bias toward determining stories are logically consistent, while Gemini-3-Pro shows the opposite tendency. As a result, although some models display high performance on either the positive set (stories with no errors) or the negative set, the overall performance is poor. Human performance is also not particularly high, largely due to the lengthy context. If we highlight the events containing the error and the relevant knowledge, their localization performance on the negative set increases from 0.61 to 0.86.
@@
Takeaway Most state-of-the-art LLMs show near-random performance in knowledge state tracking and potential action estimation. Test-time scaling does not show a clear advantage.